Ross preferences are convex, complete, and transitive. He likes two goods: Rags (R) and Paper (P). He prefers more of each good to less. His budget constraint is represented by the following equation: 2P 5R-40. Which of the following are plausible optimal bundles for Ross? Select all that apply. 8 rags, 0 papers 5 rags, 10 papers. 3 rags, 15 papens 2 rags, 15 papers
Answer #1

Options (1) and (4).

The optimal bundle will exhaust the budget. From budget line it's seen that price of P is 2 and price of R is 5.

Total cost (TC) = 2P + 5R

For option (1), TC = 2 x 0 + 5 x 8 = 0 + 40 = 40 (Exhausts budget)

For option (2), TC = 2 x 10 + 5 x 5 = 20 + 25 = 45 (Exceeds budget)

For option (3), TC = 2 x 15 + 5 x 3 = 30 + 15 = 45 (Exceeds budget)

For option (4), TC = 2 x 15 + 5 x 2 = 30 + 10 = 40 (Exhausts budget)